Prove that if az = 4; az = 7 and an = 2an-1 - An-2 +2 for n 22 then an = n2 + 3.
Answer #1

Prove that if a = 4; 9₂= 7 and an=2am-1-Am-2+2 for m2 Az = 2 a₂ - a + 2 2 (7) - 4 + 2 = 14-4+2 03 12 = (3)?+ 3 2 az-a + 2 = 2